OpenStudy (anonymous):

An arrow is shot straight up in the air at an initial speed of 15.0 m/s. After how much time is the arrow moving downward at a speed of 8.00 m/s?

OpenStudy (anonymous):

final velocity = 0 0 = 15 + (-9.8)(t) t = 1.53 s That is only to go up. Coming down, initial velocity = 0 final velocity = 8 a = 9.8 t = (8-0)/9.8 =0.816s Add them up = 2.346s

OpenStudy (anonymous):

Awesome thanks!

OpenStudy (anonymous):

Alright quick questions, when I am working out the downward time I am getting a negitive number. Does that just indicate the direction of movement like with acceleration?

OpenStudy (anonymous):

gravity! set it's direction right.

OpenStudy (anonymous):

Time will ALWAYS be positive, no matter what. You need to cancel out the -ve signs using acceleration and velocity.

OpenStudy (anonymous):

Ok so when i'm going down gravity is positive then if i'm fighting gravity its negitve correct?

OpenStudy (turingtest):

You can define either direction as positive so long as you are consistent in your work

OpenStudy (anonymous):

Yes. That is how most people learn it.

OpenStudy (anonymous):

Awesome guys thanks a ton!